Please help me solve this age problem: Rachelle is now 2 less than thrice as old as May. In 5 yrs, May is as old as Rachelle was 3 yrs ago. How old are they now?
Answer by TimothyLamb(4379)

x = rachelle age now
y = may age now
---
x = 3y - 2
x - 3 = y + 5
---
put the system of linear equations into standard form
---
x - 3y = -2
x - y = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = rachelle age now = 13
y = may age now = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php